OpenStudy (anonymous):

Given the terms a1=5 and a4=14 of an arithmetic sequence, what is a20?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

a[4] - a[1] = 14 - 5 = 9 So d = 9/3 = 3. This means we're adding 3 to each term to get the next term ------------------------------- a[n] = a[1] + (n-1)*d a[n] = 5 + (n-1)*3 a[n] = 5 + 3n-3 a[n] = 3n + 2 So the nth term is 3n+2 This means that the 20th term is 3n+2 3*20+2 60+2 62 So the 20th term is 62

jimthompson5910 (jim_thompson5910):

So a20 = 62

OpenStudy (anonymous):

Thank you sooo much! =)

jimthompson5910 (jim_thompson5910):

you're welcome again:)